### v1

@@ 3.3 Mamba with Other Architectures
To combine Mamba with convolution, Mamba introduces the ability to obtain local information, which is essential for tasks related to medical images or segmentation tasks. RES-VMAMBA chen2024res pioneers incorporating a residual learning framework within the VMamba model to simultaneously leverage global and local state features inherent in the original VMamba architectural design. This architecture commences with a stem module responsible for processing the input image, followed by a series of VSS Blocks organized sequentially across four distinct stages. Diverging from the original VMamba framework, the Res-VMamba architecture adopts the VMamba structure as its backbone and directly integrates raw data into the feature map. They refer to this integration as the global-residual mechanism to distinguish it from the residual structure in the VSS block. This integration aims to facilitate the sharing of global image features alongside the information processed through the VSS blocks. This design seeks to harness the localized details captured by individual VSS blocks and the overarching global features inherent in the unprocessed input, thereby enhancing the model’s representational capacity and improving its performance on tasks requiring a comprehensive understanding of visual data.
@@
To harness the long-sequence modeling capabilities of Mamba blocks and the spatiotemporal representation prowess of LSTM, the VMRNNtang2024vmrnn Cell eliminates all weights and biases in ConvLSTMshi2015convolutional and employs VSS blocks to learn spatial dependencies vertically. In the VMRNN Cell, long-term and short-term temporal dependencies are captured by updating the information on cell states and hidden states from a horizontal perspective. Building upon the VMRNN Cell, two variants are proposed: VMRNN-B and VMRNN-D. VMRNN-B primarily relies on stacking VMRNN layers, while VMRNN-D incorporates more VMRNN Cells and introduces Patch Merging and Patch Expanding layers. The Patch Merging layer serves for downsampling, effectively reducing the spatial dimensions of the data, which aids in decreasing computational complexity and capturing more abstract, global features. Conversely, the patch-expanding layer is utilized for upsampling, increasing the spatial dimensions to restore detail and enable precise localization of features in the reconstruction phase. Ultimately, the reconstruction layer takes the hidden state from the VMRNN layer and scales it back to the input size, generating the predicted frame for the next time step. Integrating downsampling and upsampling processes offers significant advantages in our predictive architecture. Downsampling simplifies the input representation, enabling the model to process higher-level features with reduced computational overhead. This is particularly advantageous for more abstractly understanding complex patterns and relationships within the data.
@@
The SSM-ViT block zubic2024state is introduced for effective event-based information processing. It comprises three main components: a self-attention block (Block-SA), a dilated attention block (Grid-SA), and an SSM block. Block-SA focuses on immediate spatial relations and provides a detailed representation of nearby features. Grid-SA offers a global perspective, capturing comprehensive spatial relations and overall input structure. The SSM block ensures temporal consistency and smooth information transition between consecutive time steps. By integrating SSM with self-attention, the SSM-ViT block enables faster training and parameter timescale adjustment for temporal aggregation.
The Meet More Areas (MMA) block introduced in cheng2024activating adopts a MetaFormer-style architecture, comprising two Layer Normalization layers, a token mixer (consisting of a channel attention mechanism and a ViM block in parallel), and an MLP block for deep feature extraction. There are two main reasons for this choice: Firstly, models adopting MetaFormer-style architectures have shown promising results, indicating the potential for achieving favorable outcomes. Secondly, to fully leverage and utilize the global information extracted by the ViM block, the channel attention mechanism is incorporated to activate more pixels, as global details play a role in determining the channel attention weights. Additionally, it’s reasonable to suggest that employing a convolution-based module can enhance the visual representation obtained by the ViM block and streamline the training process, similar to the benefits observed with transformers. For restoration, Residual State Space Blocks (RSSBs) guo2024mambair block add VSS block in front of the channel attention block, which enables VSS to focus on learning diverse channel representation after which the critical channels are selected by subsequent channel attention, thus avoiding channel redundancy.
@@ 4 Visual Mamba in Application Fields
Backbone Visual Mamba Vision Mamba  zhu2024vision
@@
SiMBA  patro2024simba -
RES-VMAMBA  chen2024res -
Efficient Mamba EfficientVMamba pei2024efficientvmamba -
MambaMixer behrouz2024mambamixer -
High/Mid-level vision Object detection SSM-ViT zubic2024state Params 17.5
Segmentation ReMamber  yang2024remamber -
Sigma  wan2024sigma -
Video classification ViS4mer  islam2022long Memory 5273.6
Video understanding Video Mamba Suite  chen2024video -
VideoMamba  li2024videomamba
@@
SpikeMba  li2024spikemba -
Multi-Modal understanding Cobra  zhao2024cobra -
ReMamber  yang2024remamber -
VL-Mamba  qiao2024vl -
Video prediction VMRNN  tang2024vmrnn Params 2.6, FLOPs 0.9
HARMamba li2024harmamba
@@
Low-level vision Image super-resolution MMA  cheng2024activating -
Image restoration MambaIR  guo2024mambair Params 16.7
SERPENT  shahab2024serpent -
VmambaIR  shi2024vmambair Params 10.50, FLOPs 20.5
Image dehazing UVM-Net  zheng2024u Params 19.25
Image derain FreqMamba  zhen2024freqmamba Params 14.52
Image deblurring ALGNet  gao2024aggregating FLOPs 17
Visual generation MambaTalk  xu2024mambatalk -
Motion Mamba  zhang2024motion -
DiS  fei2024scalable -
ZigMa  hu2024zigma -
Point cloud 3DMambaComplete  li20243dmambacomplete
@@
3DMambaIPF  zhou20243dmambaipf -
Point Cloud Mamba  zhang2024point
@@
POINT MAMBA  liu2024point Memory 8550
SSPointMamba  liang2024pointmamba
@@
3D reconstruction GAMBA  shen2024gamba -
Video generation SSM-based diffusion model  oshima2024ssm -
{tablenotes}
@@ 4.1.1 High/Mid-level vision
The Visual Mamba Backbone zhu2024vision; liu2024vmamba; yang2024plainmamba; huang2024localmamba; li2024mamba performance decent on classification, object detection and segmentation. SSM-ViT zubic2024state is designed for object detection using event cameras. Unlike standard frame-based cameras, event cameras record per-pixel relative brightness changes in a scene as they occur. Therefore, object detection with event cameras requires processing an asynchronous stream of events in a four-dimensional spatio-temporal space. Earlier studies have employed RNN architectures with convolutional or attention mechanisms to develop models that exhibit superior performance on downstream tasks using event camera data. However, these models often suffer from slow training. In response, the SSM-ViT block is introduced, leveraging SSM for efficient event-based information processing. It explores two strategies to mitigate aliasing effects when deploying the model at higher frequencies.
@@
Leveraging Mamba’s significant advancements in efficient training and inference with linear complexity, ReMamber yang2024remamber is introduced for referring image segmentation (RIS), a challenging task in the realm of multi-modal understanding. Distinguished from conventional segmentation, RIS entails identifying and segmenting specific objects in images based on textual descriptions. The ReMamber architecture comprises several Mamba Twister blocks, each featuring multiple VSS blocks and a Twisting layer. The Mamba Twister block serves as a multi-modal feature fusion block, receiving visual and textual features as input and outputting the fused multi-modal feature representation. Intermediate features are extracted after each Mamba Twister block and subsequently fed into a flexible decoder to generate the final segmentation mask. The VSS layers are tasked with extracting visual features, while the Twisting layer primarily captures effective visual-language interactions. The authors conduct experiments on multiple RIS datasets, achieving state-of-the-art results.Sigma wan2024sigma is a novel network for multimodal semantic segmentation tasks. One of them, Siam Mamba encoder, uses cascaded visual state space (VSS) blocks to extract multi-scale global information from different modalities. And an attention-based Mamba fusion mechanism and a channel-aware Mamba decoder are proposed. In the decoding stage, the fused features are further enhanced by channel-aware visual state space (CVSS) blocks, which can effectively capture multi-scale long-range information and realize cross-modal information integration.
@@
Unlike transformers that rely on quadratic complexity attention mechanisms, Mamba, as a pure SSM-based model, excels in handling long sequences with linear complexity and proves particularly adept at processing lengthy videos at high resolutions. ViS4mer islam2022long serves as a model primarily used for recognizing and classifying long videos, particularly for understanding and categorizing lengthy movie clips. ViS4mer consists of two main components: a standard Transformer encoder designed for extracting short-distance spatiotemporal features from videos, and a multi-scale temporal S4 decoder optimized for subsequent long-range temporal reasoning. Leveraging the capability of the fundamental SSM to capture long-range dependencies in sequential data, the multi-scale temporal S4 decoder is implemented based on SSMs to reduce the computation cost of the model.
@@
The Video Mamba Suite chen2024video is not a novel method; rather, it explores and assesses the potential of SSM, represented by Mamba, in video understanding tasks. The ViM block is enhanced into the Decomposed Bidirectionally Mamba (DBM) Block, which separates the input projector while sharing the parameters of SSM in both scanning directions. They classify Mamba into four distinct roles for modeling videos and compare it with existing Transformer-based models to evaluate its effectiveness in various video understanding tasks. Furthermore, the Video Mamba Suite comprises 14 models/modules to evaluate performance across 12 video understanding tasks. The experiments demonstrate that Mamba is applicable in video analysis and can be utilized for more complex, multimodal video understanding challenges. Apart from the Video Mamba Suite, VideoMamba li2024videomamba is proposed for video understanding tasks, with a specific focus on addressing two major challenges: local redundancy and global dependencies. This study evaluates VideoMamba’s capabilities across four key aspects: scalability in the video domain, sensitivity to short-term action recognition, advantages in long-term video understanding, and compatibility with other modalities. To enhance model scalability in the visual domain, VideoMamba employs a self-distillation strategy. This approach significantly enhances VideoMamba’s performance as both the model and input sizes increase, without the need for pre-training on large-scale datasets. While the ViM block enhances the model’s spatial perception capabilities, VideoMamba extends this capability to 3D video understanding by including spatio-temporal bidirectional scanning. Through the extension of the ViM block, VideoMamba achieves a notable increase in processing speed and a decrease in computational resource consumption without compromising performance.SpikeMba li2024spikemba is a novel multimodal video content understanding framework designed to handle the task of temporal video localization in video content.SpikeMba combines Spiking Neural Networks (SNNs) and State Space Models (SSMs) to capture fine-grained relationships between multimodal input features. In particular, the Spike Saliency Detector (SSD) utilizes the thresholding mechanism of SNNs to generate saliency proposal sets that signal highly relevant or salient instances in the video through spikes. The Multimodal Relevance Mamba Block (MRM) is based on SSM and enhances the modeling of long range dependencies while maintaining linear complexity with respect to the input size.
@@
In recent years, Multimodal Large Language Models (MLLMs) have been used with remarkable success in various domains. However, as a base model for many downstream tasks, current MLLMs consist of well-known Transformer networks with low secondary computational complexity. To improve the efficiency of these models, Cobrazhao2024cobra has been proposed. This is a new multimodal large language model (MLLM) with linear computational complexity, which integrates the efficient Mamba language model into the visual modality, exploring different modal fusion schemes, and thus identifying ways to produce the most efficient multimodal representation.Cobra consists of three components: a visual encoder, a projector, and a Mamba backbone. Among them, the visual coder is used to extract the visual representation of the image, and the projector is used to transform the dimensions of the visual representation to match the dimensions of the tokens in the Mamba language model.The Mamba backbone, on the other hand, consists of a stack of 64 identical basic blocks that receive a combination of visual and textual embeddings while preserving the connectivity and the RMSNorm, and autoregressively transforms them into the target token sequences. VL-Mamba qiao2024vl is also a multimodal large language model consisting of a pre-trained visual coder, a randomly initialized MMC, and a pre-trained Mamba LLM. Among them, the visual coder uses the Vision Transformer (ViT) architecture to generate a sequence of patch features of the original image.For the MMC, a 2D visual selective scanning mechanism is proposed to solve the computer vision task since the state-space model is designed to deal with 1D sequential data with causality and the visual sequences generated by the visual coder are 2D non-causal data . The article explores three multimodal connector variants: the MLP (multilayer perceptron), the VSS-MLP (MLP combined with a VSS module) and the VSS-L2 (two linear layers combined with a VSS module). Input images are first acquired as visual features through a visual coder, then visual sequences are fed into the MMC, and finally the output vectors are fed into the Mamba LLM in combination with a tokenized text query in order to generate the corresponding response.The integration and processing of visual and verbal information is optimized through the synergy of these components.The Referential Image Segmentation (RIS) task requires a model to recognize and segment specific objects in an image based on textual descriptions, and ReMamberyang2024remamber is a novel architecture for handling this task.ReMamber combines the Mamba model and introduces multimodal Mamba Twister blocks to explicitly simulate image-text interactions through its unique channeling and spatial warping mechanism Fusing textual and visual features.ReMamber extracts intermediate features after each Mamba Twister block and feeds them into a flexible decoder to generate the final segmentation mask.
@@
To address the unparalleled challenge of predicting temporal and spatial dynamics for spatio-temporal forecasting in videos, the VMRNN cell tang2024vmrnn introduces a novel recurrent unit designed to efficiently handle spatio-temporal prediction tasks. Recognizing the challenges in processing extensive global information, the VMRNN cell integrates VSS blocks with LSTM architecture to leverage the long-sequence modeling abilities of VSS blocks and the spatio-temporal representation capabilities of LSTM. This integration enhances the accuracy and efficiency of spatio-temporal predictions. The model conducts image-level analysis by segmenting each frame into patches, which are subsequently flattened and processed through an embedding layer. This process enables the VMRNN layer to extract and predict spatio-temporal features effectively. HARMamba li2024harmamba builds on ViT blocks for activity recognition and reaches superior performance while reducing computational and memory overhead on the activity recognition tasks.

### v2

@@ 3.3 Mamba with Other Architectures
To combine Mamba with convolution, Mamba introduces the ability to obtain local information, which is essential for tasks related to medical images or segmentation tasks. RES-VMAMBA [48] pioneers incorporating a residual learning framework within the VMamba model to simultaneously leverage global and local state features inherent in the original VMamba architectural design. This architecture commences with a stem module responsible for processing the input image, followed by a series of VSS Blocks organized sequentially across four distinct stages. Diverging from the original VMamba framework, the Res-VMamba architecture adopts the VMamba structure as its backbone and directly integrates raw data into the feature map. They refer to this integration as the global-residual mechanism to distinguish it from the residual structure in the VSS block. This integration aims to facilitate the sharing of global image features alongside the information processed through the VSS blocks. This design seeks to harness the localized details captured by individual VSS blocks and the overarching global features inherent in the unprocessed input, thereby enhancing the model’s representational capacity and improving its performance on tasks requiring a comprehensive understanding of visual data.
@@
To harness the long-sequence modeling capabilities of Mamba blocks and the spatiotemporal representation prowess of LSTM, the VMRNN[47] Cell eliminates all weights and biases in ConvLSTM[70] and employs VSS blocks to learn spatial dependencies vertically. In the VMRNN Cell, long-term and short-term temporal dependencies are captured by updating the information on cell states and hidden states from a horizontal perspective. Building upon the VMRNN Cell, two variants are proposed: VMRNN-B and VMRNN-D. VMRNN-B primarily relies on stacking VMRNN layers, while VMRNN-D incorporates more VMRNN Cells and introduces Patch Merging and Patch Expanding layers. The Patch Merging layer serves for downsampling, effectively reducing the spatial dimensions of the data, which aids in decreasing computational complexity and capturing more abstract, global features. Conversely, the patch-expanding layer is utilized for upsampling, increasing the spatial dimensions to restore detail and enable precise localization of features in the reconstruction phase. Ultimately, the reconstruction layer takes the hidden state from the VMRNN layer and scales it back to the input size, generating the predicted frame for the next time step. Integrating downsampling and upsampling processes offers significant advantages in our predictive architecture. Downsampling simplifies the input representation, enabling the model to process higher-level features with reduced computational overhead. This is particularly advantageous for more abstractly understanding complex patterns and relationships within the data.
@@
The SSM-ViT block [71] is introduced for effective event-based information processing. It comprises three main components: a self-attention block (Block-SA), a dilated attention block (Grid-SA), and an SSM block. Block-SA focuses on immediate spatial relations and provides a detailed representation of nearby features. Grid-SA offers a global perspective, capturing comprehensive spatial relations and overall input structure. The SSM block ensures temporal consistency and smooth information transition between consecutive time steps. By integrating SSM with self-attention, the SSM-ViT block enables faster training and parameter timescale adjustment for temporal aggregation.
The Meet More Areas (MMA) block introduced in [41] adopts a MetaFormer-style architecture, comprising two Layer Normalization layers, a token mixer (consisting of a channel attention mechanism and a ViM block in parallel), and an MLP block for deep feature extraction. There are two main reasons for this choice: Firstly, models adopting MetaFormer-style architectures have shown promising results, indicating the potential for achieving favorable outcomes. Secondly, to fully leverage and utilize the global information extracted by the ViM block, the channel attention mechanism is incorporated to activate more pixels, as global details play a role in determining the channel attention weights. Additionally, it’s reasonable to suggest that employing a convolution-based module can enhance the visual representation obtained by the ViM block and streamline the training process, similar to the benefits observed with transformers. For restoration, Residual State Space Blocks (RSSBs) [57] block add VSS block in front of the channel attention block, which enables VSS to focus on learning diverse channel representation after which the critical channels are selected by subsequent channel attention, thus avoiding channel redundancy.
@@ 4 Visual Mamba in Application Fields
Backbone Visual Mamba Vision Mamba  [26]
@@
SiMBA  [33] -
RES-VMAMBA  [48] -
Efficient Mamba EfficientVMamba [30] -
MambaMixer [31] -
High/Mid-level vision Object detection SSM-ViT [71] Params 17.5
Segmentation ReMamber  [50] -
Sigma  [49] -
Video classification ViS4mer  [72] Memory 5273.6
Video understanding Video Mamba Suite  [43] -
VideoMamba  [36]
@@
SpikeMba  [73] -
Multi-Modal understanding Cobra  [74] -
ReMamber  [50] -
VL-Mamba  [42] -
Video prediction VMRNN  [47] Params 2.6, FLOPs 0.9
HARMamba [40]
@@
Low-level vision Image super-resolution MMA  [41] -
Image restoration MambaIR  [57] Params 16.7
SERPENT  [58] -
VmambaIR  [35] Params 10.50, FLOPs 20.5
Image dehazing UVM-Net  [75] Params 19.25
Image derain FreqMamba  [68] Params 14.52
Image deblurring ALGNet  [76] FLOPs 17
Visual generation MambaTalk  [77] -
Motion Mamba  [37] -
DiS  [78] -
ZigMa  [34] -
Point cloud 3DMambaComplete  [79]
@@
3DMambaIPF  [80] -
Point Cloud Mamba  [81]
@@
POINT MAMBA  [44] Memory 8550
SSPointMamba  [82]
@@
3D reconstruction GAMBA  [83] -
Video generation SSM-based diffusion model  [84] -
  • 1
@@ 4.1.1 High/Mid-level vision
The Visual Mamba Backbone [26, 27, 29, 28, 32] performance decent on classification, object detection and segmentation. SSM-ViT [71] is designed for object detection using event cameras. Unlike standard frame-based cameras, event cameras record per-pixel relative brightness changes in a scene as they occur. Therefore, object detection with event cameras requires processing an asynchronous stream of events in a four-dimensional spatio-temporal space. Earlier studies have employed RNN architectures with convolutional or attention mechanisms to develop models that exhibit superior performance on downstream tasks using event camera data. However, these models often suffer from slow training. In response, the SSM-ViT block is introduced, leveraging SSM for efficient event-based information processing. It explores two strategies to mitigate aliasing effects when deploying the model at higher frequencies.
@@
Leveraging Mamba’s significant advancements in efficient training and inference with linear complexity, ReMamber [50] is introduced for referring image segmentation (RIS), a challenging task in the realm of multi-modal understanding. Distinguished from conventional segmentation, RIS entails identifying and segmenting specific objects in images based on textual descriptions. The ReMamber architecture comprises several Mamba Twister blocks, each featuring multiple VSS blocks and a Twisting layer. The Mamba Twister block serves as a multi-modal feature fusion block, receiving visual and textual features as input and outputting the fused multi-modal feature representation. Intermediate features are extracted after each Mamba Twister block and subsequently fed into a flexible decoder to generate the final segmentation mask. The VSS layers are tasked with extracting visual features, while the Twisting layer primarily captures effective visual-language interactions. The authors conduct experiments on multiple RIS datasets, achieving state-of-the-art results.Sigma [49] is a novel network for multimodal semantic segmentation tasks. One of them, Siam Mamba encoder, uses cascaded visual state space (VSS) blocks to extract multi-scale global information from different modalities. And an attention-based Mamba fusion mechanism and a channel-aware Mamba decoder are proposed. In the decoding stage, the fused features are further enhanced by channel-aware visual state space (CVSS) blocks, which can effectively capture multi-scale long-range information and realize cross-modal information integration.
@@
Unlike transformers that rely on quadratic complexity attention mechanisms, Mamba, as a pure SSM-based model, excels in handling long sequences with linear complexity and proves particularly adept at processing lengthy videos at high resolutions. ViS4mer [72] serves as a model primarily used for recognizing and classifying long videos, particularly for understanding and categorizing lengthy movie clips. ViS4mer consists of two main components: a standard Transformer encoder designed for extracting short-distance spatiotemporal features from videos, and a multi-scale temporal S4 decoder optimized for subsequent long-range temporal reasoning. Leveraging the capability of the fundamental SSM to capture long-range dependencies in sequential data, the multi-scale temporal S4 decoder is implemented based on SSMs to reduce the computation cost of the model.
@@
The Video Mamba Suite [43] is not a novel method; rather, it explores and assesses the potential of SSM, represented by Mamba, in video understanding tasks. The ViM block is enhanced into the Decomposed Bidirectionally Mamba (DBM) Block, which separates the input projector while sharing the parameters of SSM in both scanning directions. They classify Mamba into four distinct roles for modeling videos and compare it with existing Transformer-based models to evaluate its effectiveness in various video understanding tasks. Furthermore, the Video Mamba Suite comprises 14 models/modules to evaluate performance across 12 video understanding tasks. The experiments demonstrate that Mamba is applicable in video analysis and can be utilized for more complex, multimodal video understanding challenges. Apart from the Video Mamba Suite, VideoMamba [36] is proposed for video understanding tasks, with a specific focus on addressing two major challenges: local redundancy and global dependencies. This study evaluates VideoMamba’s capabilities across four key aspects: scalability in the video domain, sensitivity to short-term action recognition, advantages in long-term video understanding, and compatibility with other modalities. To enhance model scalability in the visual domain, VideoMamba employs a self-distillation strategy. This approach significantly enhances VideoMamba’s performance as both the model and input sizes increase, without the need for pre-training on large-scale datasets. While the ViM block enhances the model’s spatial perception capabilities, VideoMamba extends this capability to 3D video understanding by including spatio-temporal bidirectional scanning. Through the extension of the ViM block, VideoMamba achieves a notable increase in processing speed and a decrease in computational resource consumption without compromising performance.SpikeMba [73] is a novel multimodal video content understanding framework designed to handle the task of temporal video localization in video content.SpikeMba combines Spiking Neural Networks (SNNs) and State Space Models (SSMs) to capture fine-grained relationships between multimodal input features. In particular, the Spike Saliency Detector (SSD) utilizes the thresholding mechanism of SNNs to generate saliency proposal sets that signal highly relevant or salient instances in the video through spikes. The Multimodal Relevance Mamba Block (MRM) is based on SSM and enhances the modeling of long range dependencies while maintaining linear complexity with respect to the input size.
@@
In recent years, Multimodal Large Language Models (MLLMs) have been used with remarkable success in various domains. However, as a base model for many downstream tasks, current MLLMs consist of well-known Transformer networks with low secondary computational complexity. To improve the efficiency of these models, Cobra[74] has been proposed. This is a new multimodal large language model (MLLM) with linear computational complexity, which integrates the efficient Mamba language model into the visual modality, exploring different modal fusion schemes, and thus identifying ways to produce the most efficient multimodal representation.Cobra consists of three components: a visual encoder, a projector, and a Mamba backbone. Among them, the visual coder is used to extract the visual representation of the image, and the projector is used to transform the dimensions of the visual representation to match the dimensions of the tokens in the Mamba language model.The Mamba backbone, on the other hand, consists of a stack of 64 identical basic blocks that receive a combination of visual and textual embeddings while preserving the connectivity and the RMSNorm, and autoregressively transforms them into the target token sequences. VL-Mamba [42] is also a multimodal large language model consisting of a pre-trained visual coder, a randomly initialized MMC, and a pre-trained Mamba LLM. Among them, the visual coder uses the Vision Transformer (ViT) architecture to generate a sequence of patch features of the original image.For the MMC, a 2D visual selective scanning mechanism is proposed to solve the computer vision task since the state-space model is designed to deal with 1D sequential data with causality and the visual sequences generated by the visual coder are 2D non-causal data . The article explores three multimodal connector variants: the MLP (multilayer perceptron), the VSS-MLP (MLP combined with a VSS module) and the VSS-L2 (two linear layers combined with a VSS module). Input images are first acquired as visual features through a visual coder, then visual sequences are fed into the MMC, and finally the output vectors are fed into the Mamba LLM in combination with a tokenized text query in order to generate the corresponding response.The integration and processing of visual and verbal information is optimized through the synergy of these components.The Referential Image Segmentation (RIS) task requires a model to recognize and segment specific objects in an image based on textual descriptions, and ReMamber[50] is a novel architecture for handling this task.ReMamber combines the Mamba model and introduces multimodal Mamba Twister blocks to explicitly simulate image-text interactions through its unique channeling and spatial warping mechanism Fusing textual and visual features.ReMamber extracts intermediate features after each Mamba Twister block and feeds them into a flexible decoder to generate the final segmentation mask.
@@
To address the unparalleled challenge of predicting temporal and spatial dynamics for spatio-temporal forecasting in videos, the VMRNN cell [47] introduces a novel recurrent unit designed to efficiently handle spatio-temporal prediction tasks. Recognizing the challenges in processing extensive global information, the VMRNN cell integrates VSS blocks with LSTM architecture to leverage the long-sequence modeling abilities of VSS blocks and the spatio-temporal representation capabilities of LSTM. This integration enhances the accuracy and efficiency of spatio-temporal predictions. The model conducts image-level analysis by segmenting each frame into patches, which are subsequently flattened and processed through an embedding layer. This process enables the VMRNN layer to extract and predict spatio-temporal features effectively. HARMamba [40] builds on ViT blocks for activity recognition and reaches superior performance while reducing computational and memory overhead on the activity recognition tasks.